OpenStudy (imamermaid29):

WILL fan & medal The system of equations is inconsistent. What are the missing values? 5x + ___y = 2 ___x + 3y = 8

OpenStudy (anonymous):

Okay. An easy way to determine if two lines are parallel from the given form above... Ax + By = E Cx + Dy = F If the constants are a multiple of one another than the lines are parallel. If you can multiply, Ax+By by some number , and get Cx+Dy, the lines will be parallel.

OpenStudy (anonymous):

o if you let the first equation be... 5x + 1y = 2 you can see that the second equation will be 3 times the first, The first will have 1y, and the second has 3y. So you have to multiply the 5x by 3

OpenStudy (anonymous):

3*(5x + 1y) ---> 15x + 3y Here is a good reference for these types of probs... http://www.algebra.com/algebra/homework/coordinate/Types-of-systems-inconsistent-dependent-independent.lesson

OpenStudy (anonymous):

A step further, if you multiply the first equation by a constant number, and not only the values on the X and Y become the same, but also the value on the Right of the = sign becomes the same, then Both are the same line. For example 4X + 5Y = 5 8X + 10Y =10 The second equation is just 2 times the first equation. Both are the same line, they overlap. They have Infinitely many solutions, and are dependent.

OpenStudy (anonymous):

Hope this helped! Have a great day! @ImaMermaid29